Given that "A^{-1}=\frac{1}{\textup{det}(A)}\begin{bmatrix}d&-b\\-c&a\end{bmatrix}"  with "\textup{det}(A)\neq 0", consider the following statements:

(i) "\frac{1}{\textup{det}(A)}\begin{bmatrix}d&-b\\-c&a\end{bmatrix}=\frac{1}{\textup{det}(A)}\textup{adj}(A)" ;


(ii) "A=\begin{bmatrix}a&c\\b&d\end{bmatrix}^T" ;


(iii) "\textup{adj}(A)=\begin{bmatrix}d&-c\\-b&a\end{bmatrix}^T"


Which of the above are true?



1
Expert's answer
2021-09-14T06:09:41-0400

We know that for an invertible matrix A

"A^{-1}=\frac{1}{det(A)}.adJ(A)....(I)"

For a 2×2 matrix "A=\begin{bmatrix}
 a & b \\
 c & d
\end{bmatrix}"

"A^{-1}=\frac{1}{det(A)}\begin{bmatrix}
 d & -b \\
 -c & a
\end{bmatrix}......(2)"

We gave

"A^{-1}=\frac{1}{det(A)}\begin{bmatrix}
 d & -b \\
 -c & a
\end{bmatrix}\space with A\#0"


Therefore from equition 1 we can conclude that

"A^{-1}=\frac{1}{det(A)}.adJ(A)=\frac{1}{det(A)}\begin{bmatrix}
 d & -b \\
 -c & a
\end{bmatrix}.....(2)"

Since we have

"A^{-1}=\frac{1}{det(A)}\begin{bmatrix}
 d & -b \\
 -c & a
\end{bmatrix}"


Therefore

"A=\begin{bmatrix}
 a & b \\
 c & d
\end{bmatrix}"

"A=\begin{bmatrix}
 a & b \\
 c & d
\end{bmatrix}^{T}....(4)"

Also

adj(A)"=\begin{bmatrix}
 d & -b \\
 -c & a
\end{bmatrix}=\begin{bmatrix}
 d & -c \\
 -b & a
\end{bmatrix}^{T}....(5)"


Hence, proved that (I)(ii) and (iii) are all true.